OpenStudy (anonymous):

A man is on a diet and goes into a shop to buy some turkey slices. he is given 3 slices which together weigh 1/3 of a pound, but his diet allows 1/4 of a pound. how much of the 3 slices can he eat while staying true to his diet? I need help explaining this and how to solve this equation!

OpenStudy (anonymous):

3 slices = 1/3 of a pound implies 9 slices = 1 pound (I just multiplied both sides by 3) But we want to find out how many slices is equal to 1/4 of a pound, so lets divide by 4: 9/4 slices = 1/4 pound 9/4 = 2.25 slices!

OpenStudy (anonymous):

oooh I see!! wow much better understanding! thank you so much!